bon wants to take an early retirement in 10 years when he will be twice as old as his daughter Sonia. Sonia is 25 years younger than her father. How old is bon now? hoe old will he be when he retired? (please use one variable)(thanks in advance)
Answer by ikleyn(52776)

.

The solution using two variables and two equations

Let x be the Bon's age and let y be the Sonia's (his daughter's) age.


From one side, we have this equation 

x - y = 25            (1)    ("Sonia is 25 years younger than her father".)



From the other side, we have second equation

x + 10 = 2*(y+10)     (2)    (in 10 years . . . )


So, we have the system of equations (1) and (2) in two unknowns.

To solve it, express  x = 25 + y  from eq(1), and substitute it into equation (2). You will get

(25 + y) + 10 = 2*(y+10)

35 + y = 2y + 20

35 - 20 = 2y - y

15 = y.


Thus Sonia is 15 years old now.

Hence, the father is  x = 25 + y = 25 + 15 = 40 years old.


Answer.  Sonia is 15 years old;  the father is 40 years old.

Solved.


The solution using one variable and single equation

Let x be the Sonia's age.

Then the Bon's age is  (x+25), according to the condition.


In 10 years Sonia's age will be (x+10) years,  while Bon's age will be (x+25)+10 = (x+35).


Then the condition says

x + 35 = 2*(x+10).


It implies

x + 35 = 2x + 20

35 - 20 = 2x - x

15 = x.


And you get THE SAME ANSWER.

The second solution is completed.

------------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.